xx CarpetIOBasic

CarpetIOBasic

Erik Schnetter <schnetter@cct.lsu.edu>

December 20 2005

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11intxwidth 10Scope: privatex01 INT

11Description: Field width for integer values

11  Range  10  x01  Default: 9

111:*  10

11

11iterxwidth  10Scope: privatex01  INT

11Description: Field width for the current iteration

11  Range  10  x01  Default: 9

111:*  10

11

11outheaderxevery  10Scope: privatex01  INT

11Description: How often to print the header

11 Range 10 x01 Default: 20
111:*  10Output every so many time steps
11-1  10No header output

11

11outinfoxcriterion  10Scope: privatex01  KEYWORD

11Description: Criterion to select scalar output intervals, overrides outxevery

11 Range 10 x01 Default: iteration
11never  10Never output
11iteration  10Output every so many iterations
11divisor  10Output if iteration mod divisor == 0.
11time  10Output every that much coordinate time

11

11outinfoxdt  10Scope: privatex01  REAL

11Description: How often to do scalar output, overrides outxdt

11 Range 10 x01 Default: -2
11(0:*  10In intervals of that much coordinate time
11                   10As often as possible
11-1  10Disable output
11-2  10Default to IO::outxdt

11

11outinfoxevery  10Scope: privatex01  INT

11Description: How often to do scalar output, overrides IO::outxevery

11 Range 10 x01 Default: -2
111:*  10Output every so many time steps
11-1:0  10No output
11-2  10Default to IO::outxevery

11

11outinfoxreductions 10Scope: privatex01  STRING

11Description: List of reductions to output in scalar form

11 Range 10 x01 Default: minimum maximum
11                   10A regex which matches everything

11

11outinfoxvars  10Scope: privatex01  STRING

11Description: Variables to output in scalar form

11 Range 10 x01 Default: (none)
11                   10A regex which matches everything

11

11realxmax  10Scope: privatex01  REAL

11Description: Upper bound for numbers that are displayed in fixed notation

11 Range 10 x01 Default: 1.0e+3
11(0.0:*  10

11realxmin  10Scope: privatex01  REAL

11Description: Lower bound for numbers that are displayed in fixed notation

11 Range 10 x01 Default: 1.0e-8
11(0.0:*  10

11

11realxprec  10Scope: privatex01  INT

11Description: Precision for real values

11  Range  10  x01  Default: 7

110:*  10

11

11realxprecxsci  10Scope: privatex01  INT

11Description: Precision for real values in scientific notation

11  Range  10  x01  Default: 6

110:*  10

11

11realxwidth  10Scope: privatex01  INT

11Description: Field width for real values

11  Range  10  x01  Default: 12

111:*  10

11

11timexprec  10Scope: privatex01  INT

11Description: Precision for the simulation time

11  Range  10  x01  Default: 3

110:*  10

11

11timexwidth  10Scope: privatex01  INT

11Description: Field width for the simulation time

11  Range  10  x01  Default: 9

111:*  10

11

11outxcriterion  10Scope: shared from IOx01KEYWORD

11

11outxdt  10Scope: shared from IOx01REAL

11

11outxevery  10Scope: shared from IOx01INT

11

11strictxioxparameterxcheck 10Scope: shared from IOx01BOOLEAN

11

7 Interfaces General Implements: iobasic

Inherits: io

Grid Variables

7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11lastxoutputxiteration10 lastxoutputxiteration x01  compact110
11                  10  x01 dimensions110
11                  10  x01 distribution11CONSTANT
11                  10  x01 group type11SCALAR
11                  10  x01 timelevels111
11                  10  x01variable type11INT
11 lastxoutputxtime 10 lastxoutputxtime x01  compact110
11                  10  x01 dimensions110
11                  10  x01 distribution11CONSTANT
11                  10  x01 group type11SCALAR
11                  10  x01 timelevels111
11                  10  x01variable type11REAL
11 thisxiteration 10 thisxiteration x01  compact110
11                  10  x01 dimensions110
11                  10  x01 distribution11CONSTANT
11                  10  x01 group type11SCALAR
11                  10  x01 timelevels111
11                  10  x01variable type11INT

11

Uses header: Timer.hh carpet.hh typecase.hh typeprops.hh

8 Schedule This section lists all the variables which are assigned storage by thorn Carpet/CarpetIOBasic. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11x
11lastxoutputxiteration lastxoutputxtime thisxiteration11x

11x  11x

11

Scheduled Functions

CCTKxSTARTUP x carpetiobasicstartup x startup routine x

11x11After: 11ioutilxstartup
11 x 11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function

11

CCTKxINITIAL x carpetiobasicinit x initialisation routine x

11x11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function
11 x 11Writes: 11carpetiobasic::lastxoutputxiteration
11 x 11x  11carpetiobasic::lastxoutputxtime
11 x 11x  11carpetiobasic::thisxiteration